OpenStudy (anonymous):

You need to move a 105kg sofa to a diffrent room. It takes a force of 102 N to start it moving. What is the coefficient of static friction between sofa and carpet?

OpenStudy (anonymous):

for \[\sum Fy=ma \] =N-mg=0 N=105 * 9.81 =???? then \[\sum Fx=ma\] F-f =0 \[F=uN\] \[ 102N=(105*9.81)u\] \[u=???? \]